OpenStudy (anonymous):

Marge drove 348 miles in four and a half hours. At that rate, how many miles could she drive in 10 hours?

hero (hero):

348/4.5 = ? What ever the answer is to that, multiply it by 10

hero (hero):

348/4.5 = 77.3333 77.3333 * 10 = 773.3333 miles